OpenStudy (unklerhaukus):

A person of mass \(m\) jumps off the ground. Suppose the person pushes off the ground with a constant force of magnitude \(F\) for \(T\) seconds. What was the magnitude of the displacement of the centre of mass of the person while they were in contact with the ground?

OpenStudy (experimentx):

is information complete?

OpenStudy (unklerhaukus):

Express your answer in terms of m, F, T, and g as needed.

OpenStudy (anonymous):

8.01x??

OpenStudy (anonymous):

0.5*(F/(m)-g)*T^2

OpenStudy (anonymous):

use the motion equations. s=ut+0.5at^2 where s=distance traveled by center of mass Rcm u=0= initial velocity a= acceleration comes from force calculation i.e. force down=mg force up=F Therefore acceleration=(F-mg)/m Substitute in the equation at the top Rcm=0+0.5*((F/m)-g)*T^2